OpenStudy (anonymous):

Solve for x and I need to see the steps x^2e^-x-xe^-x-12e^-x=0

OpenStudy (lalaly):

factor out e^-x\[e^{-x}(x^2-x-12)=0\]

OpenStudy (lalaly):

\[x^2-x-12=0\]

OpenStudy (lalaly):

\[(x-4)(x+3)=0\]

OpenStudy (anonymous):

Let e^-x be b bx^2 - bx - 12b = 0 b (x^2 - x - 12) = 0 b ( x + 3) ( x - 4) = 0 x = -3, 4